OpenStudy (anonymous):

can you simplify sqrt e^2+81if so to what?

OpenStudy (anonymous):

the square root of 81 is 9... n the square root of e^2 is e... other than that, i don't think can simplify it further

OpenStudy (anonymous):

is it sqrt of e^(2+81) or sqrt of (e^2)+81

OpenStudy (anonymous):

it is \[\sqrt{e^2+81}\]

OpenStudy (anonymous):

no that cant be siplified

OpenStudy (anonymous):

i thought than it would be (e+9)

OpenStudy (anonymous):

no, cause (e+9)^2=e^2+18e+81

OpenStudy (anonymous):

that is true so no you can't simplify this than

OpenStudy (anonymous):

thus, its not the square root

OpenStudy (anonymous):

oh yeah... great thinking pumpkin!

OpenStudy (anonymous):

ok now Im confused

OpenStudy (anonymous):

no, you cannot simplify

OpenStudy (anonymous):

that is the simplest form

OpenStudy (anonymous):

ok thanks guys your awsome